Circular No. 7230

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

PROBABLE NEW SATELLITES OF URANUS J. J. Kavelaars, McMaster University; B. Gladman, Observatoire de la

Côte d’Azur (OCA); M. Holman, Harvard-Smithsonian Center for Astrophysics (CfA); J.-M. Petit, OCA; and H. Scholl, OCA, report the discovery of two probable new satellites of Uranus in CCD images obtained with the 3.5-m Canada-France-Hawaii Telescope at Mauna Kea:

1999 UT  α_{2000}  δ_{2000}  R
July 18.51250  21^{h}14^{m}55.^{s}58  −16^{o}39′31′′.9  23.4  S/1999 U 1
18.59355  21 14 54.87  −16 39 35.1
19.50227  21 14 47.10  −16 40 12.5
19.57032  21 14 46.47  −16 40 15.3
21.46464  21 14 29.75  −16 41 35.2  23.1
21.51919  21 14 29.26  −16 41 37.5
18.51250  21 13 38.33  −16 49 57.1  24.3  S/1999 U 2
18.59355  21 13 37.61  −16 50 00.6
19.50227  21 13 29.54  −16 50 39.5
19.57032  21 13 28.89  −16 50 42.4
21.46464  21 13 11.76  −16 52 06.7  24.3
21.51919  21 13 11.26  −16 52 09.3

Although computations show that the data can be satisfied equally well by satellite or by centaur orbits, the absence of other slow-moving objects nearby with apparent motions remotely resembling that of Uranus suggests that the objects are indeed satellites. The ephemerides below are from computations by B. G. Marsden, CfA, in which it was assumed that the objects are near the apocenters of direct uranicentric orbits having e = 0.4:

1999 TT α_{2000} δ_{2000} Δ r_{U} ϵ Δα′′ Δδ′′
July 21 21^{h}14.^{m}57 −16^{o}41.^{′}2 18.959 0.154 161.^{o}9 1544+ 625+
31 21 13.05 −16 48.4 18.922 0.154 171.8 1557+ 612+
Aug. 10 21 11.47 −16 55.7 18.913 0.154 178.1 1567+ 597+
20 21 09.90 −17 02.9 18.934 0.154 168.2 1573+ 581+
30 21 08.40 −17 09.6 18.984 0.154 158.2 1576+ 564+
July 21 21 13.27 −16 51.7 18.889 0.062 162.2 425+  5−
31 21 11.71 −16 59.3 18.853 0.061 172.1 402+ 40−
Aug. 10 21 10.09 −17 06.9 18.846 0.061 177.7 377+ 74−
20 21 08.48 −17 14.4 18.870 0.060 167.8 349+ 107−
30 21 06.94 −17 21.4 18.923 0.060 157.8 318+ 139−
1999 July 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Brian G. Marsden